# Some insights into depth estimators for location and scatter in the multivariate setting

### v1

@@ 5.1 Maximum bias of the deepest scatter estimator
Remark 18
$\Phi\left(l_{1}^{1/2}\right)$ and $\Phi\left(l_{p}^{1/2}\right)$ are the quantities predicted by the bounds in Chen et al (2018), $B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right)-1\right]$ and $B_{I}\left(\varepsilon\right)=1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right).$ Both functions goes to the the maximum and minimum value when $\varepsilon=1/3.$
Lemma 19
Let us consider the case of having a family of matrices $\Gamma_{r}$ with the largest eigenvalue $l_{1}^{(r)}\rightarrow$infty, $r^{2}l_{1}^{{}^{(r)}-1/2}\rightarrow 1/2,$ and the other eigenvalues going to 0. Then, $\lim_{r\rightarrow\infty}D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left(\varepsilon,1-\varepsilon\right).$
Proof. We have to analyze,
$\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{r}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\leq\frac{1}{\sum_{j=2}^{p}l_{j}^{{}^{(r)}-1}\left(\mathbf{w}_{r}^{t}\mathbf{e}_{j}\right)^{2}}\right)+\\ \varepsilon\delta\left(r^{2}l_{1}^{{}^{(r)}-1/2}\mathbf{w}_{r}^{t}\mathbf{ee}^{t}\mathbf{w}_{r}\leq 1\right),\\ \left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{r}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\geq\frac{1}{\sum_{j=2}^{p}l_{j}^{{}^{(r)}-1}\left(\mathbf{w}_{r}^{t}\mathbf{e}_{j}\right)^{2}}\right)+\\ \varepsilon\delta\left(r^{2}l_{1}^{{}^{(r)}-1/2}\mathbf{w}_{r}^{t}\mathbf{ee}^{t}\mathbf{w}_{r}\geq 1\right)\end{array}\right\}$
@@
(i) If $\ \mathbf{w}_{r}$ $\in L\left(\mathbf{e}_{2},\dots,\mathbf{e}_{p}\right)$ we get for $r$ sufficiently large,
$\min\left\{\varepsilon,\left(1-\varepsilon\right)\right\}.$
(ii) If $\mathbf{w}_{r}$ $\in\left[L\left(\mathbf{e}_{2},\dots,\mathbf{e}_{p}\right)\right]^{c},$ we get for $r$ sufficiently large,
$\min\left\{\varepsilon,\left(1-\varepsilon\right)\right\}.$
Then, for $r$ large enough, since $\varepsilon\leq 1/2,$
$D\left(\Gamma_{r},P_{\varepsilon,r}\right)=\varepsilon.$ (18)
and we conclude the statement of the lemma.
Lemma 20
The critical points of $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)$ (respectively $1-\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right))$ subject to $\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\geq 0$ (idem $\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\leq 0)$ are either $\mathbf{v}_{p},$ $\mathbf{v}_{1}.$ or occurs at $F_{r}^{\mathbf{e}}.$
Proof. Take the Lagrangian $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)+\lambda\left(\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\right)+\gamma\left(\mathbf{v}^{t}\mathbf{v-}1\right),$ Karush, Kuhn, Tucker condition says that if the minimum local occurs in $\mathbf{\tilde{v},}$ then either $\mathbf{\tilde{v}}^{t}\Gamma\mathbf{\tilde{v}-}r^{2}\left(\mathbf{\tilde{v}}^{t}\mathbf{e}\right)^{2}=0$ or we have to consider the Lagrangian $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)+\gamma\left(\mathbf{v}^{t}\mathbf{v-}1\right).$ In the last case, it was shown that the critical points are the eigenvectors of $\Gamma.$ Suppose that $\mathbf{v}_{j_{0}\text{,}}$ $j_{0}<p,$is the eigenvector corresponding to the minimum $l_{j_{0}}$ such that $\mathbf{v}_{j_{0}\text{ }}^{t}\Gamma\mathbf{v}_{j_{0}\text{ }}>r^{2}\left(\mathbf{v}_{j_{0}\text{ }}^{t}\mathbf{e}\right)^{2}$, by the continuity of the functions we could find another $\mathbf{\bar{v}}$ such that $\mathbf{v}_{j_{0}\text{ }}^{t}\Gamma\mathbf{v}_{j_{0}\text{ }}>\mathbf{\bar{v}}_{\text{ }}^{t}\Gamma\mathbf{\bar{v}}>r^{2}\left(\mathbf{\bar{v}}^{t}\mathbf{e}\right)^{2}>r^{2}\left(\mathbf{v}_{j_{0}\text{ }}^{t}\mathbf{e}\right)^{2}$  which is a contradiction. Therefore the critical points occur at $\mathbf{\tilde{v}=v}_{p}$ $\left(\mathbf{v}_{1}\right)$ or $\mathbf{\tilde{v}}^{t}\Gamma\mathbf{\tilde{v}-}r^{2}\left(\mathbf{\tilde{v}}^{t}\mathbf{e}\right)^{2}=0.$
The following lemma will give us the maximum bias for the deepest estimator under point mass contaminations.
Lemma 21
If $\hat{\Gamma}$ stands for a deepest estimator, then .$D\left(\hat{\Gamma},P_{\varepsilon,r}\right)\leq\left(1-\varepsilon\right)/2$.
Proof. Given $\Gamma=\sum l_{v}\mathbf{v}_{j}\mathbf{v}_{j}^{t}$.The depth of a matrix is given by, if $r>l_{1}^{1/2}.$
@@
$\begin{array}[]{cccccccccc}A_{r}^{\mathbf{e}}&\mathbf{v}_{1},\mathbf{v}_{p}&&&\mathbf{v}_{1}&\mathbf{v}_{p}&\mathbf{v}_{1}&\mathbf{v}_{p}&&\\ B_{r}^{\mathbf{e}}&&\mathbf{v}_{1},\mathbf{v}_{p}&&\mathbf{v}_{p}&\mathbf{v}_{1}&&&\mathbf{v}_{p}&\mathbf{v}_{1}\\ F_{r}^{\mathbf{e}}&&&\mathbf{v}_{1},\mathbf{v}_{p}&&&\mathbf{v}_{p}&\mathbf{v}_{1}&\mathbf{v}_{1}&\mathbf{v}_{p}\end{array}$
@@
In any case, we have that $D\left(\Gamma,P_{\varepsilon,r}\right)\leq\min\left\{\left(1-\varepsilon\right)g_{s},\left(1-\varepsilon\right)\left(1-g_{t}\right)\right\},s\in\left\{p,m\right\},$ $t\in\left\{1,M\right\},$ $g_{s}\leq 1-g_{t}$ and the statement follows.
Theorem 22
The asymptotic breakdown point of the deepest estimator is $1/3$
Proof. (i) Lemma above says that if the estimators moves to the boundary of the domain of the eigenvalues, then $\varepsilon\geq 1/3$  and $\varepsilon^{\ast}\geq 1/3.$
(ii) If the depth estimators $\left\{\hat{\Gamma}_{r}\right\}$ remain bounded in the $\varepsilon-$neighborhood, we have that for point mass contaminatios we have that $D\left(\Gamma_{r},P_{\varepsilon,r}\right)\leq D\left(\hat{\Gamma}_{r},P_{\varepsilon,r}\right)$ for $r$ large enough. By the preceding lemmas, we have that $\varepsilon\leq\left(1-\varepsilon\right)/2,$ and $\varepsilon\leq 1/3.$Then, if the depth estimator remains bounded in the $\varepsilon-$contamination neighborhood then the level of contamination $\varepsilon$ is less than $1/3$ Equivalently, if $\varepsilon>1/3$ then the depth estimator becomes unbounded in the $\varepsilon-$contamination neighborhood. Therefore $\varepsilon^{\ast}\leq 1/3.$
Next, we follow closely to Theorem 4.1 and 4.2 of Chen and Tyler (2002). Let us take the contaminated distribution $P_{\varepsilon,Q}=\left(1-\varepsilon\right)P_{0}+\varepsilon Q$ and $A\subset\left\{\Gamma\succeq\mathbf{0}\right\}$, definir
Lemma 23
  1. 1.
  2. 2.
  3. 3.
    $B\left(\hat{\Gamma},\varepsilon,P_{0}\right)\leq\left\|L\left(\frac{\varepsilon}{1-\varepsilon},P_{0}\right)\right\|=\max\left\{\frac{1}{\beta}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),\frac{\beta}{\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)}\right\}$
Proof. Since $P_{0}=\Phi$ then $D_{M}\left(\Phi\right)=1/2$ Observe that $\delta\geq 0$ since given $Q$ any distribution on $\mathbb{R}^{p},$
$P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$$\geq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\text{ \ },\text{ \ \ }$
$P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$$\geq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\geq$ $\left(1-\varepsilon\right)\min\left\{P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
@@
$D\left(\Gamma,P_{\varepsilon,Q}\right)$ $=$ $\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
@@
$\min\left\{\begin{array}[]{c}P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$$\leq$ $P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\leq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)+\varepsilon$
$\min\left\{\begin{array}[]{c}P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$$\leq$ $P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\leq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)+\varepsilon$
@@
$\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\leq$ $\left(1-\varepsilon\right)\min\left\{P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}+\varepsilon$
@@
$D\left(\Gamma,P_{\varepsilon,Q}\right)\leq\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)+\varepsilon$
and (i) follows.
(ii) Set $\alpha=\frac{e}{1-\varepsilon}-\delta\left(\varepsilon,P_{0}\right).$ Then $20\leq\alpha<1/$ 2 .  If $\Gamma\notin L\left(\alpha.P_{0}\right)$ we have that $D\left(\Gamma,P_{0}\right)<D_{M}\left(P_{0}\right)-\alpha.$ Therefore,
@@
(iii) $L\left(\alpha,P_{0}\right)=L\left(\varepsilon/\left(1-\varepsilon\right)-\delta,P_{0}\right)\subset.L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ Then, $\Gamma\in\left(L\left(\alpha,P_{0}\right)\right)^{c}$ implies, by  (ii) that $\Gamma\in\left(M\left(P_{\varepsilon,Q}\right)\right)^{c}$ or $\left(L\left(\alpha,P_{0}\right)\right)^{c}\subset\left(M\left(P_{\varepsilon,Q}\right)\right)^{c}$ which says that $M\left(P_{\varepsilon,Q}\right)\subset L\left(\alpha,P_{0}\right)\subset L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ for all distribution $Q.$
Since $D\left(\Gamma,P_{0}\right)=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right)$ and $D_{M}\left(P_{0}\right)=1/2,$
@@
Then, $\left\|\Gamma\right\|_{op}\leq\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right)$ and $\left\|\Gamma^{-1}\right\|_{op}\leq 1/\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)$ if $\Gamma\in L\left(\varepsilon/\left(1-\varepsilon\right),P\right)$ and

### v2

@@ 7.1 Maximum bias of the deepest scatter estimator
[…]
Proof of Lemma 4.9 We have to analyze,
$\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\mathbf{XX}^{t}\frac{\Gamma_{r}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\leq\frac{1}{\sum_{j=2}^{p}l_{j}^{{}^{(r)}-1}\left(\mathbf{w}_{r}^{t}\mathbf{e}_{j}\right)^{2}}\right)+\\ \varepsilon\delta\left(r^{2}l_{1}^{{}^{(r)}-1/2}\mathbf{w}_{r}^{t}\mathbf{ee}^{t}\mathbf{w}_{r}\leq 1\right),\\ \left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\mathbf{XX}^{t}\frac{\Gamma_{r}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\geq\frac{1}{\sum_{j=2}^{p}l_{j}^{{}^{(r)}-1}\left(\mathbf{w}_{r}^{t}\mathbf{e}_{j}\right)^{2}}\right)+\\ \varepsilon\delta\left(r^{2}l_{1}^{{}^{(r)}-1/2}\mathbf{w}_{r}^{t}\mathbf{ee}^{t}\mathbf{w}_{r}\geq 1\right)\end{array}\right\},$
@@
[…]
(ii) If $\mathbf{w}_{r}$ $\in\left[L\left(\mathbf{e}_{2},\dots,\mathbf{e}_{p}\right)\right]^{c},$ we get for $r$ sufficiently large, $\min\left\{\varepsilon,\left(1-\varepsilon\right)\right\}.$ Then, for $r$ large enough, since $\varepsilon\leq 1/2,$
$D\left(\Gamma_{r},P_{\varepsilon,r}\right)=\varepsilon,$ (31)
[…]
Proof of Lemma 4.10. Take $\Gamma=\sum_{j=1}^{p}l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t}$. The depth of a matrix is given by, if $r>l_{1}^{1/2},$
@@
$A_{r}^{\mathbf{e}}$ $\mathbf{v}_{1},\mathbf{v}_{p}$ $\mathbf{v}_{1}$ $\mathbf{v}_{p}$ $\mathbf{v}_{1}$ $\mathbf{v}_{p}$
$B_{r}^{\mathbf{e}}$ $\mathbf{v}_{1},\mathbf{v}_{p}$ $\mathbf{v}_{p}$ $\mathbf{v}_{1}$ $\mathbf{v}_{p}$ $\mathbf{v}_{1}$
$F_{r}^{\mathbf{e}}$ $\mathbf{v}_{1},\mathbf{v}_{p}$ $\mathbf{v}_{p}$ $\mathbf{v}_{1}$ $\mathbf{v}_{1}$ $\mathbf{v}_{p}$
@@
[…]
Proof of Theorem 4.11. (i) Lemma above says that if the estimators moves to the boundary of the domain of the eigenvalues, then $\varepsilon\geq 1/3$ and $\varepsilon^{\ast}\geq 1/3.$
[…]
Proof of Lemma 4.12. Since $P_{0}=\Phi$ then $D_{M}\left(\Phi\right)=1/2$ . Observe that $\delta\geq 0$ since given $Q$ any distribution on $\mathbb{R}^{p},$
[…]
$P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$ $\geq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\geq$ $\left(1-\varepsilon\right)\min\left\{P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
@@
$D\left(\Gamma,P_{\varepsilon,Q}\right)$ $=$ $\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
@@
$\min\left\{\begin{array}[]{c}P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$ $\leq$ $P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\leq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)+\varepsilon$
$\min\left\{\begin{array}[]{c}P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$ $\leq$ $P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\leq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)+\varepsilon$
@@
$\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\leq$ $\left(1-\varepsilon\right)\min\left\{P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}+\varepsilon$
@@
$D\left(\Gamma,P_{\varepsilon,Q}\right)\leq\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)+\varepsilon.$
Consequently, (i) follows.
(ii) Set $\alpha=\frac{e}{1-\varepsilon}-\delta\left(\varepsilon,P_{0}\right).$ Then $0\leq\alpha<1/2.$ If $\Gamma\notin L\left(\alpha.P_{0}\right)$ we have that $D\left(\Gamma,P_{0}\right)<D_{M}\left(P_{0}\right)-\alpha.$ Therefore,
@@
(iii) $L\left(\alpha,P_{0}\right)=L\left(\varepsilon/\left(1-\varepsilon\right)-\delta,P_{0}\right)\subset L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ Then, $\Gamma\in\left(L\left(\alpha,P_{0}\right)\right)^{c}$ implies, by (ii) that $\Gamma\in\left(M\left(P_{\varepsilon,Q}\right)\right)^{c}$ or $\left(L\left(\alpha,P_{0}\right)\right)^{c}\subset\left(M\left(P_{\varepsilon,Q}\right)\right)^{c}$ which says that $M\left(P_{\varepsilon,Q}\right)\subset L\left(\alpha,P_{0}\right)\subset L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ for all distribution $Q.$ Since $D\left(\Gamma,P_{0}\right)=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right)$ and $D_{M}\left(P_{0}\right)=1/2,$ we have that
@@
Then, $\left\|\Gamma\right\|_{op}\leq\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),$ $\left\|\Gamma^{-1}\right\|_{op}\leq 1/\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)$ if $\Gamma\in L\left(\varepsilon/\left(1-\varepsilon\right),P\right)$ and